OpenStudy (anonymous):

What value of x makes the ratios equivalent? thirty over five equals eighteen over x

OpenStudy (anonymous):

30/5 = 18/x

OpenStudy (texaschic101):

30/5 = 18/x cross multiply (30)(x) = (18)(5) 30x = 90 x = 90/30 x = 3 check... 30/5 = 18/x 30/5 = 18/3 cross multiply (30)(3) = (18)(5) 90 = 90 (correct) ANSWER: x = 3